OpenStudy (anonymous):

what is the geometric mean of 125 and 7

OpenStudy (anonymous):

Let a=125 and b=7 The the geometric mean of a and b is : \[\sqrt{ab}=\sqrt{125\times7}=\sqrt{875}\]

OpenStudy (anonymous):

the choices are a)55 b)45 c)35 d)40 @Noura11 ??

OpenStudy (anonymous):

no chance this is a whole number

OpenStudy (anonymous):

its 175 and 7 my bad.... @Noura11 @satellite73

OpenStudy (anonymous):

then you know what to do right?

OpenStudy (amistre64):

does 7 divide 175?

OpenStudy (johnweldon1993):

Then you would still use that same equation above...just with 175 and 7 \[\sqrt{ab}=\sqrt{175\times7}=\sqrt{1225}\]